Write and Equation in Slope Intercept Form for the line containing the points (3,-17) and (-4,18)
Answer by TimothyLamb(4379)

slope = dy/dx
m = (18+17)/(-4-3)
m = -5
---
y - 18 = -5(x + 4)
y - 18 = -5x - 20
y = -5x - 20 + 18
y = -5x - 2
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---